OpenStudy (anonymous):

sketch a graph of y=f(x) f(x)=-3x+2

OpenStudy (zzr0ck3r):

what is the slope? what is the y intercept?

OpenStudy (anonymous):

take note that the equation is in the form: y = mx +b, where m = slope and b = y-intercept.

OpenStudy (anonymous):

example: f(x) = -4x + 1 is also y = -4x +1. Thus, slope = -4, and y-intercept = 1. Which means we have a point (0,1).

OpenStudy (anonymous):

to graph the equation on the example, plot the (0,1) (point representing the y-intercept). then, consider that slope = rise/run = -4/1. So, 4 units downward, then 1 unit to the right.

OpenStudy (anonymous):

 then, connect the two points by a line.